OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 30 miles per gallon of gas and the second has a fuel efficiency of 25 miles per gallon of gas. During one particular week, the two cars went a combined total of 2050 miles, for a total gas consumption of 75 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (anonymous):

I think it is 37.5 gallons

OpenStudy (mathstudent55):

Car A does 30 mpg and used x gallons of gas. Car B does 25 mpg and used y gallons of gas. We are told that the total gas used is 75 gallons, so we have x + y = 75 If we multiply the gas mileage by the gallons of each car, we find the miles each car traveled. The total miles traveled is 2050. 30 * x + 25 * y = 2050 We need to solve the system of equation: x + y = 75 30x + 25y = 2050